OpenStudy (anonymous):

How do you find the marginal cost of C(x)= 4x R(x)= 8-0.001x^2

OpenStudy (anonymous):

dC/dx=4=marginal cost

OpenStudy (anonymous):

I thought it was just the derivative but I am confused since there are 2 equations..... I also have to find marginal revenue and marginal profit so maybe that is why there is a 2nd equation??

OpenStudy (anonymous):

Oh so it is just the derivative. How would you find marginal revenue then please?

OpenStudy (anonymous):

dR/dx=-0.002x=marginal revenue

OpenStudy (anonymous):

but I guess R and C should changed to each other.

OpenStudy (anonymous):

meant C=8-0.001x2 R=4x

OpenStudy (anonymous):

The original equation was 8x-0.001x^2 but I see. So then the marginal profit would be what?

OpenStudy (anonymous):

Pi=R-C dR/dx=dC/dx->8-0.002x=4->x=4/0.002=2000

OpenStudy (anonymous):

Oh ok I understand. Thank you!!